OpenStudy (anonymous):

How much energy is required to move the electron sufficiently far away from the proton such that it does not experience the proton's electric field. Express your answer in units of Joules.

Parth (parthkohli):

Do you know the formula for energy in Bohr's Model?

OpenStudy (anonymous):

Ya....

Parth (parthkohli):

\[E = -\dfrac{Z^2}{n^2} 13.6 ~\rm J\]

Parth (parthkohli):

You want to excite the electron to \(0 \rm J\)

OpenStudy (anonymous):

But, shouldn't I use the value of bohr radius for the initial distance between the electron and proton ?

Parth (parthkohli):

You can.

OpenStudy (anonymous):

And what value of Z should I use? the question doesn't say anything about the atomic number or the element? Should I use Z=1 as in Hydrogen?

Parth (parthkohli):

You want the electron and proton to not be affected by each other's electric field. This is possible only at infinity. If the radius in \(\infty \), then the energy is \(0~\rm J\). So you can directly apply the formula for energy.

Parth (parthkohli):

I think you should leave \(\rm Z\) as it is.

Parth (parthkohli):

\[\Delta E = E_f - E_i = \dfrac{Z^2}{n^2}\left(13.6\right) ~\rm J\]

OpenStudy (anonymous):

Ok, I think I got it! thanx for your help!

Parth (parthkohli):

Just for confirmation, what answer does it give in your book?

Parth (parthkohli):

@Sourav12084 Oh my God... lol, I gave you the answer in eV. Convert eV to Joules.

Parth (parthkohli):

Multiply that expression by \(1.6 \times 10^{-19}\)

OpenStudy (anonymous):

The question wasn't from a book...my teacher gave it for homework...and yea...i converted the answer in joules already...thanx a lot for your help...